OpenStudy (ajanijones):

What is the probability of the event when a number cube with faces labeled 1–6 is rolled once? a 2 or a 3 A. B. C. D.

OpenStudy (ajanijones):

A = 1/36 B = 1/6 C = 1/3 D = 1/2

OpenStudy (anonymous):

1/6 for the number 2, 1/6 for the number 3 2 chances out of 6 to get a 2 or a 3 C = 2/6 = 1/3